# UniAP: Unifying Inter- and Intra-Layer Automatic Parallelism by Mixed Integer Quadratic Programming

### v5

@@ 1 Introduction
Deep learning models have demonstrated promising performance across many domains. For example, deep learning models such as BERT [5], GPT-3 [1] and T5 [31] achieve state-of-the-art (SOTA) performance on many natural language processing (NLP) tasks. For computer vision (CV), deep learning models such as ViT [6] and Swin Transformer [24] achieve good accuracy on multiple tasks.
@@ 2.1 Parallel strategy
Pipeline parallelism (PP) In PP, each worker (machine or GPU) holds a subset of model layers. Adjacent layers on different workers need to transfer activations in the forward propagation (FP) step and gradients in the backward propagation (BP) step.
Data parallelism (DP) In DP, each worker holds a replica of the whole model and partitions training samples. In each iteration, each worker computes gradients and synchronizes them with the other workers using all-reduce collective communication (CC). All workers will have the same model parameters after the synchronization step.
Tensor parallelism (TP) In TP, each worker holds a replica of training samples and partitions within model layers. In each iteration, each worker computes its local outputs in FP and its local gradients in BP. To synchronize outputs and gradients, all workers will perform all-reduce CC in FP and BP steps according to the partition scheme.
Fully sharded data parallelism (FSDP) FSDP partitions optimizer states, parameters and gradients of the model into separate workers. During the FP and BP step of each iteration, FSDP performs an all-gather CC to obtain the complete parameters for the relevant layer, respectively. After computing gradients, FSDP conducts a reduce-scatter CC to distribute the global gradients among the workers.
@@ 2.2 Parallel method
(a) Model
(b) MP
(c) Inter-layer-only AP
(d) Intra-layer-only AP
(e) Hierarchical AP
(f) UniAP
Figure 1: Parallel methods for optimizing parallel strategies for a three-layer model in (a). Numbered boxes represent the corresponding layers of the model. Layers with gray backgrounds represent the final parallel strategy for each method. Layers without dotted lines indicate that DP was applied to them, while TP was applied to the others. Solid arrows between layers represent connections within a single pipeline stage, while dashed arrows represent connections between consecutive pipeline stages.
@@ 3 Method
In this section, we introduce our proposed method called UniAP, which jointly optimizes the two categories of parallel strategies, including PP, DP, TP, and FSDP, to find an optimal solution. Figure 1 illustrates the difference between UniAP and other parallel methods. MP methods are manually designed with limited flexibility. Inter-layer-only and intra-layer-only AP methods optimize (search) from a set of candidate inter-layer-only and intra-layer-only parallel strategies, respectively. Hierarchical AP methods first adopt greedy or dynamic programming to propose candidate inter-layer parallel strategies. Then, they optimize the intra-layer parallel strategy for every fixed inter-layer parallel strategy. UniAP has the largest strategy space for exploration (joint optimization).
@@ 3.3.2 Constraint
Detailed proof can be found in Appendix A.
@@ 3.4 Unified optimization process
First, UOP optimizes intra-layer-only parallelism for cases in which pipeline parallelism is not adopted. Several works [47, 23] have used quadratic integer programming (QIP) to optimize intra-layer-only parallel strategy and achieved promising results. UniAP provides a QIP formulation for intra-layer-only parallelism in Appendix B.
@@
For each candidate $deg$ and $c$, UOP formulates the cost for a training iteration to an MIQP expression. It then waits for the MIQP solver to return the optimal cost and parallel strategy. Specifically, our implementation adopts the Gurobi Optimizer [13] as our QIP and MIQP solver.
Finally, UOP returns the minimum cost $cost^{*}$ and its corresponding pipeline parallel size $deg^{*}$, number of micro-batches $c^{*}$, layer placement $\textbf{{P}}^{*}$, and intra-layer strategies $\textbf{{S}}^{*}$. We provide visualization for a candidate solution to UOP in Appendix C.
@@ 4.1 Experiment setup
We conduct experiments on four different kinds of environments to validate the effectiveness and universality of our method. EnvA refers to a node (machine) with 1 Xeon 6248 CPU, 8 V100-SXM2 32GB GPUs, and 472GB memory. EnvB refers to 4 nodes interconnected with 10Gbps networks and each node has 2 Xeon E5-2620 v4 CPUs, 4 TITAN Xp 12GB GPUs, and 125GB memory. EnvC refers to a node with 8 A100 40GB PCIe GPUs. EnvD is a cloud cluster with 16 nodes and each node has 4 DCUs (a type of non-NVIDIA GPU). In the following text, we will specify the number of GPUs used after the environment name (e.g., EnvB-8) only when partial nodes of the environment are used.
We evaluate UniAP with five Transformer-based models, BERT-Huge [5], T5-Large [31], ViT-Huge [6], Swin-Huge [24], and Llama [40, 41]. We follow the common practice of training these transformer-based models. To eliminate factors that affect training throughput, we turn off techniques orthogonal to parallel strategies, such as activation checkpointing [3]. However, we integrate FP16 mixed precision training [26] for the largest model, Llama, in some cases to successfully orchestrate the process. More details on these models are provided in Appendix D.
The experimental evaluation concentrates on two primary metrics: training throughput and strategy optimization time. The former is calculated by averaging throughput from the 10th to the 60th iteration of training, while the latter is determined by measuring the time of the UOP. More details are provided in Appendix D. Besides, we provide the accuracy of our cost model in Appendix E.
@@ 4.2 Training throughput and strategy optimization time
On EnvB-8, UniAP consistently demonstrates competitive or larger training throughput compared to Galvatron and Alpa. We attribute the performance improvement to the larger strategy space in UniAP. A detailed study is provided in Section 4.4. Furthermore, the strategy optimization time of UniAP is also significantly shorter than the two baseline methods.
@@
To facilitate further discussions, we provide a case study of the optimal parallel strategy in Appendix F.
@@ 5 Conclusion
Appendix A Proof of the linear form for the contiguous set
@@
Theorem A.1.
A subgraph $\mathbb{V}_{i}=\{\forall u\in\mathbb{V}:\textbf{{P}}_{ui}=1\}$ is contiguous if and only if there exists $\textbf{{Z}}_{vi}$ such that Equation (6a), (6b), and (6c) are satisfied.
@@
Appendix B QIP formulation for intra-layer-only parallelism
@@
Appendix C Visualization for the candidate solution
@@
Appendix D Experiment detail
@@
Appendix E Estimation accuracy
@@
Appendix F Case study: BERT-Huge

### v6

@@ 2.1 Parallel strategy
Pipeline parallelism (PP) In PP [15], each worker (machine or GPU) holds a subset of model layers. Adjacent layers on different workers need to transfer activations in the forward propagation (FP) step and gradients in the backward propagation (BP) step.
Data parallelism (DP) In DP [22], each worker holds a replica of the whole model and partitions training samples. In each iteration, each worker computes gradients and synchronizes them with the other workers using all-reduce collective communication (CC). All workers will have the same model parameters after the synchronization step.
Tensor parallelism (TP) In TP [30], each worker holds a replica of training samples and partitions within model layers. In each iteration, each worker computes its local outputs in FP and its local gradients in BP. To synchronize outputs and gradients, all workers will perform all-reduce CC in FP and BP steps according to the partition scheme.
Fully sharded data parallelism (FSDP) FSDP [32, 8] partitions optimizer states, parameters and gradients of the model into separate workers. During the FP and BP step of each iteration, FSDP performs an all-gather CC to obtain the complete parameters for the relevant layer, respectively. After computing gradients, FSDP conducts a reduce-scatter CC to distribute the global gradients among the workers.
@@ 2.2 Parallel method
Figure 1: Parallel methods for optimizing parallel strategies for a three-layer model. The different arrangements of slices with varying transparency within the same layer block indicate different intra-layer parallelism strategies adopted by layers. The different arrangements of gray blocks which wrap the layer blocks indicate different inter-layer parallelism strategies.
@@ 3 Method
In this section, we introduce our proposed method called UniAP, which jointly optimizes the two categories of parallel strategies, including PP, DP, TP, and FSDP, to find an optimal solution. Figure 1 illustrates the difference between UniAP and other automatic parallelism methods. Inter-layer-only and intra-layer-only AP methods optimize (search) from a set of candidate inter-layer-only and intra-layer-only parallel strategies, respectively. Hierarchical AP methods first adopt greedy or dynamic programming to propose candidate inter-layer parallel strategies. Then, they optimize the intra-layer parallel strategy for every fixed inter-layer parallel strategy. UniAP has the largest strategy space for exploration (joint optimization).
@@ 3.3.2 Constraint
Here, Z is an auxiliary variable shaped like P. Given the node set $\mathbb{V}_{i}$ that contains nodes in stage $i$, we define $\textbf{{Z}}_{vi}=1$ if a node $w\in\mathbb{V}_{i}$ is reachable from $v~{}(v\in\mathbb{V})$. Otherwise, $\textbf{{Z}}_{vi}=0$. Detailed proof can be found in Appendix B.
@@ 3.4 Unified optimization process
First, UOP optimizes intra-layer-only parallelism for cases in which pipeline parallelism is not adopted. Several works [47, 23] have used quadratic integer programming (QIP) to optimize intra-layer-only parallel strategy and achieved promising results. UniAP provides a QIP formulation for intra-layer-only parallelism in Appendix C.
@@
For each candidate $deg$ and $c$, UOP formulates the cost for a training iteration to an MIQP expression. It then waits for the MIQP solver to return the optimal cost and parallel strategy.
Finally, UOP returns the minimum cost $cost^{*}$ and its corresponding pipeline parallel size $deg^{*}$, number of micro-batches $c^{*}$, layer placement $\textbf{{P}}^{*}$, and intra-layer strategies $\textbf{{S}}^{*}$. We provide visualization for a candidate solution to UOP in Appendix D.
@@ 4.1 Experiment setup
We conduct experiments on four different kinds of environments to validate the effectiveness and universality of our method. EnvA refers to a node (machine) with 1 Xeon 6248 CPU, 8 V100-SXM2 32GB GPUs, and 472GB memory. EnvB refers to 4 nodes interconnected with 10Gbps networks and each node has 2 Xeon E5-2620 v4 CPUs, 4 TITAN Xp 12GB GPUs, and 125GB memory. EnvC refers to a node with 8 A100 40GB PCIe GPUs. EnvD is a cloud cluster with 16 nodes and each node has 4 Hygon DCUs (a type of non-NVIDIA GPU). In the following text, we specify the number of GPUs after the environment name (e.g., EnvB-8) only when partial environment nodes are used.
We evaluate UniAP with five Transformer-based models, BERT-Huge [5], T5-Large [31], ViT-Huge [6], Swin-Huge [24], and Llama [40, 41]. We follow the common practice of training these transformer-based models. To eliminate factors that affect training throughput, we turn off techniques orthogonal to parallel strategies, such as activation checkpointing [3]. However, we integrate FP16 mixed precision training [26] for the largest model, Llama, in some cases to successfully orchestrate the process. More details on these models are provided in Appendix E.
Several solvers can be used for QIP and MIQP optimization. In our implementation, we choose Gurobi [13], with detailed configurations provided in Appendix E. Exploring other solvers is left for future work.
The experimental evaluation concentrates on two primary metrics: training throughput and strategy optimization time. The former is calculated by averaging throughput from the 10th to the 60th iteration of training, while the latter is determined by measuring the time of the UOP. More details are provided in Appendix E. Besides, we provide the accuracy of our cost model in Appendix F.
@@ 4.2 Training throughput and strategy optimization time
On EnvB-8, UniAP consistently demonstrates competitive or larger training throughput compared to Galvatron and Alpa. We attribute the performance improvement to UniAP’s larger strategy space. A detailed study is provided in Section 4.4. Furthermore, UniAP’s strategy optimization time is also significantly shorter than the two baseline methods.
@@
To facilitate further discussions, we provide a case study of the optimal parallel strategy in Appendix G.
@@ 5 Conclusion
6 Acknowledgement
This work is supported by National Key R&D Program of China (No.2020YFA0713900), NSFC Project (No.12326615), the Key Major Project of the Pengcheng Laboratory (No.PCL2024A06).
@@ 6 Acknowledgement
Appendix A Notation table
Table 5 summarizes the main notations used throughout the paper.
Symbol Description
$m_{a}$ the activation memory cost
$m_{s}$ the memory cost of model states
$ps$ parameter size
$ts$ TP size
$fs$ FSDP size
$c_{dtype}$ a constant dependent on mixed precision choice
$m_{c}$ context memory cost
$deg$ the number of computation stages
$c$ the number of micro-batches
$tpi_{gpipe}$ TPI in GPipe
$m$ the memory limit for each device
$B$ mini-batch size
$n$ the number of GPUs
$fp_{i}$ forward computation time for computation stage $i$
$bp_{i}$ backward computation time for computation stage $i$
$fo_{j}$ forward communication time for communication stage $j$
$bo_{j}$ backward communication time for communication stage $j$
$p_{i}$ the cost for computation stage $i$
$o_{j}$ the cost for communication stage $j$
$\textbf{{M}}_{uk}$ the memory cost for the $k$-th intra-layer strategy of layer $u$ on a single device
$\textbf{{A}}_{uk}$ execution cost for the $k$-th intra-layer strategy of layer $u$
$\textbf{{S}}_{uk}$ whether the $k$-th parallel strategy is selected for the layer $u$
$\textbf{{P}}_{ui}$ whether layer $u$ is placed on the $i$-th computation stage
$\textbf{{R}}_{uv}$ resharding cost between layers $u$ and $v$ when they are located within the same pipeline stage
$\textbf{{R}}^{\prime}_{uv}$ resharding cost between layers $u$ and $v$ when they are located across consecutive stages
Z the auxiliary variable used for order-preserving constraint
$\mathbb{P}$ the set of cost for computation stages, which is $\{p_{1},p_{2},\dots,p_{deg}\}$
$\mathbb{O}$ the set of cost for communication stages, which is $\{o_{1},o_{2},\dots,o_{deg-1}\}$
$\mathbb{S}_{u}$ the set of intra-layer parallel strategies for layer $u$
$\mathcal{G}(\mathbb{V},\mathbb{E})$ the computation graph for the model
Table 5: Summary of the main notations.
Appendix B Proof of the linear form for the contiguous set
@@
Theorem B.1.
A subgraph with node set $\mathbb{V}_{i}=\{\forall u\in\mathbb{V}:\textbf{{P}}_{ui}=1\}$ is contiguous if and only if there exists $\textbf{{Z}}_{vi}$ such that Equation (6a), (6b), and (6c) are satisfied.
@@
Appendix C QIP formulation for intra-layer-only parallelism
@@
Appendix D Visualization for the candidate solution
@@
Appendix E Experiment detail
@@
Appendix F Estimation accuracy
@@
Appendix G Case study: BERT-Huge
@@
Appendix H Limitation
UniAP is currently designed and tested on homogeneous clusters, but incorporating automatic parallelism for training deep models on heterogeneous clusters (e.g., a cluster equipped with both NVIDIA GPUs and DCUs) is another important research topic. Given that current parallel techniques primarily target homogeneous clusters with limited emphasis on heterogeneous clusters, we have chosen to leave this topic for future exploration.